# Finite Difference Solution Ansatz approach in Least-Squares Monte Carlo

### v7

@@ 2.2.3 Error Analysis
$e^{PV,P}_{k}\leq|\mathcal{B}^{P}_{k}|+\mathcal{Y}^{P}_{k}e^{PV,P}_{k+1},$ (2.29)
where $\mathcal{Y}^{P}_{k}=\operatorname{\mathbb{E}}(\Theta(P_{k}[V^{P}_{k+1}]-Z_{k}))$, and $\leq$’ stems from a different sign of the covariance term compared to $\operatorname{\mathbb{E}}(V_{k}^{P}-V_{k})$ in Eqn.2.28, as analysed previously.
For $|\mathcal{B}^{P}_{k}|:=\operatorname{\mathbb{E}}((Z_{k}-F_{k})[\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(Z_{k}-F_{k})])$, we have, by applying Jensen’s inequality,
$|\mathcal{B}^{P}_{k}|^{2}$$\leq$ $\operatorname{\mathbb{E}}(|Z_{k}-F_{k}|^{2}|\Theta(F_{k}-Z_{k})\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})\Theta(Z_{k}-F_{k})|^{2})$
$=$ $\operatorname{\mathbb{E}}(|Z_{k}-F_{k}|^{2}\Theta(|P_{k}[V^{P}_{k+1}]-F_{k}|-|F_{k}-Z_{k}|))$
$\leq$ $\operatorname{\mathbb{E}}(|P_{k}[V^{P}_{k+1}]-F_{k}|^{2})\operatorname{\mathbb{E}}(\Theta(|P_{k}[V^{P}_{k+1}]-F_{k}|-|F_{k}-Z_{k}|))$
$=$ $\operatorname{\mathbb{E}}(|P_{k}[V^{P}_{k+1}]-F_{k}|^{2})\mathcal{X}^{P}_{k},$
where we have introduced $\mathcal{X}^{P}_{k}:=\operatorname{\mathbb{E}}(\Theta(|P_{k}[V^{P}_{k+1}]-F_{k}|-|F_{k}-Z_{k}|))$ in the penultimate line, which in practice should be small as $P_{k}[V^{P}_{k+1}]$ is close to $F_{k}$.
@@
where in the last line, the last term is exactly the local regression least-squares error with upper bound $\Xi_{k}^{P}[V_{k+1}]$ due to the projection on $P_{k}[V_{k+1}]$, which doesn’t consider accumulated error yet (as opposed to $P_{k}[V^{P}_{k+1}]$). Then we arrive at
$|\mathcal{B}^{P}_{k}|^{2}\leq\mathcal{X}^{P}_{k}\cdot((e^{PV,P}_{k+1})^{2}+\Xi_{k}^{P}[V_{k+1}]),$ (2.30)
where we have assumed during error accumulation, $\operatorname{\mathrm{Var}}(V^{P}_{k+1}-V_{k+1})\ll(e^{PV,P}_{k+1})^{2}$ and thus the former variance term is neglected.
@@
Combining above with Eqn.2.29, we have the dynamical programming solution on the upper bound of $e^{PV,P}_{k}$, denoted by $\epsilon^{PV,P}_{k}$, for a given projection $P$, to be
$\epsilon^{PV,P}_{k}=\sqrt{\mathcal{X}^{P}_{k}\cdot((\epsilon^{PV,P}_{k+1})^{2}+\Xi_{k}^{P}[V_{k+1}])}+\mathcal{Y}^{P}_{k}\epsilon^{PV,P}_{k+1},$ (2.31)
@@
with initial condition $\epsilon^{PV,P}_{M}=0$. This defines the relation between accumulated errors and local regression errors on option prices under least-squares Monte Carlo at each step, although looking simple enough but no closed-form solution for further simplification.
As we can see from Eqn.2.25, $\Xi_{k}^{P^{\text{FD-LSM}}}[V_{k+1}]$ can be much lower than $\Xi_{k}^{P^{\text{LSM}}}[V_{k+1}]$, depending on the correlation between the original and FD auxiliary problems. We expect the same order of magnitude in terms of reduction in $\epsilon^{P^{\text{FD-LSM}}}_{k}$ compared to its LSM counterpart by solving the dynamical programming problem in Eqn.2.31.
To gain more insights from Eqn.2.31, let’s assume the same local regression errors along the time direction, i.e., $\Xi_{k}^{P^{\text{LSM(FD-LSM)}}}[V_{k+1}]=\Xi^{P^{\text{LSM(FD-LSM)}}}[V]$ for all $k$. To further simplify, we put $\mathcal{X}^{P}_{k}=\mathcal{X}$ and $\mathcal{Y}^{P}_{k}=\mathcal{Y}$ for all $k$ in both LSM and FD-LSM.
Using $\tilde{\rho}(V,\hat{V}^{1D})=90\%$, $\Xi^{P^{\text{LSM}}}[V]=0.05^{2}$ and thus $\Xi^{P^{\text{FD-LSM}}}[V]=(0.05\times\sqrt{1-90\%^{2}})^{2}=0.0218^{2}$ as a numerical example along with $\mathcal{X}^{P}=0.01$ and $\mathcal{Y}^{P}=0.8$, we have the plots for (upper bound) accumulated option price errors $\epsilon^{PV,P}_{k}$ going backward as depicted in Fig.2.1.
Figure 2.1: The upper bound of accumulated option price error for LSM and FD-LSM at each step going backward. We have assumed $\Xi^{P^{\text{LSM}}}[V]=0.05^{2}$, $\tilde{\rho}(V,\hat{V}^{1D})=90\%$, $\mathcal{X}^{P}=0.01$ and $\mathcal{Y}^{P}=0.8$. The $x$-axis is $M-k$ where M=60.
It can be shown that the error is accumulated, and saturates at
$\lim_{k\to 0}\epsilon^{PV,P}_{k}=\sqrt{\frac{\Xi^{P}[V]}{\frac{(1-\mathcal{Y}^{P})^{2}}{\mathcal{X}^{P}}-1}},$ (2.32)
as going backward. Given Eqn.2.25, the upper bound of this limiting accumulated option price error in FD-LSM is effectively reduced compared to LSM, as
@@
$\lim_{k\to 0}\epsilon^{PV,P^{\text{FD-LSM}}}_{k}=\sqrt{1-\tilde{\rho}^{2}(V,\hat{V}^{1D})}\lim_{k\to 0}\epsilon^{PV,P^{\text{LSM}}}_{k}.$ (2.33)
@@
Our quantitative result in Eqn.2.32 is consistent with previous qualitative study on the same matter by Stentoft, (2014): due to the fact that the projection approximation in LSM is only utilized to estimate stopping time rather than value function, the magnitude of accumulated error is relatively small. In fact, as stated by Tsitsiklis and Van Roy, (2001) in their Theorem 1, the accumulated error (upper bound) on option value scales as $\sqrt{M}$ when $k\to 0$ by applying the projection approximation to the value function itself. Now in addition to avoiding the divergence of error accumulation under $M\to\infty$, with a customized ansatz strongly correlated to the original problem, we show that theoretically in Eqn.2.33, FD-LSM could achieve even much lower level of accumulated errors.
Error analysis in CVA: Multi-bucket setting with WWR
@@
Now we turn to the error analysis of the CVA problem in a similar manner. As one could expect, the derivation of the proof for CVA problem in Eqn.2.11 is much less involved, due to the absence of a dynamic programming control coupled with least-squares projection, i.e., the local regression error due to $\tilde{F}_{k}\to\tilde{F}^{P}_{k}$ at each time step is decoupled and independent from other steps without being accumulated.
@@
by differentiating between $\Phi^{CVA}_{k}$ in Eqn.2.12 and its proxy defined as
@@
where $h_{cpty}(\tilde{F}^{P}_{k},t_{k})$ (taking into account WWR) is approximated due to its dependence on $\tilde{F}^{P}_{k}$ subjected to the least-squares approximation.
Given the explicit functional dependence on $\tilde{F}_{i}$ ($i\in[1,k]$) in Eqn.2.35, the upper bound of $e^{CVA,P}_{k}$, denoted by $\epsilon^{CVA,P}_{k}$, can be derived as
$\begin{split}e^{CVA,P}_{k}&\leq\operatorname{\mathbb{E}}(|\Phi^{CVA}_{k}-\Phi^{CVA,P}_{k}|)\\ &\approx\operatorname{\mathbb{E}}[|\sum^{k}_{i=1}\frac{\partial\Phi^{CVA}_{k}}{\partial\tilde{F}_{i}}\cdot(\tilde{F}^{P}_{i}-\tilde{F}_{i})|]\\ &\leq\sum^{k}_{i=1}\operatorname{\mathbb{E}}[|\frac{\partial\Phi^{CVA}_{k}}{\partial\tilde{F}_{i}}\cdot(\tilde{F}^{P}_{i}-\tilde{F}_{i})|]\\ &\leq\sum^{k}_{i=1}\sqrt{\operatorname{\mathbb{E}}[\left(\frac{\partial\Phi^{CVA}_{k}}{\partial\tilde{F}_{i}}\right)^{2}]}\cdot\sqrt{\operatorname{\mathbb{E}}[(\tilde{F}^{P}_{i}-\tilde{F}_{i})^{2}]},\\ \epsilon^{CVA,P}_{k}&:=\sum^{k}_{i=1}\sqrt{\operatorname{\mathbb{E}}[\left(\frac{\partial\Phi^{CVA}_{k}}{\partial\tilde{F}_{i}}\right)^{2}]}\cdot\sqrt{\Xi^{P}_{i}[\tilde{V}]},\\ \Xi^{P}_{i}[\tilde{V}]&:=\operatorname{\mathbb{E}}[(\tilde{F}^{P}_{i}-\tilde{F}_{i})^{2}]\end{split}$ (2.36)
@@
Without loss of generality, assuming $\Xi^{P}_{i}[\tilde{V}]=\Xi^{P}[\tilde{V}]$ for all $i$ and combining above with Lemma 1, we have
$\epsilon^{CVA,P^{\text{FD-LSM}}}_{k}=\sqrt{1-\tilde{\rho}^{2}(\tilde{V},\hat{\tilde{V}}^{1D})}\epsilon^{CVA,P^{\text{LSM}}}_{k},$ (2.37)
where $\tilde{\rho}(\tilde{V},\hat{\tilde{V}}^{1D})$ is the correlation between European options $\tilde{V}$ and $\hat{\tilde{V}}^{1D}$, a direct analogue of $\tilde{\rho}(V,\hat{V}^{1D})$ in the previous problem. It is worth to note that the CVA error under LSM could be difficult to control, since value function is approximated directly; in the new method, by virtual of the 1D ansatz as a control variate, it provides an invaluable error reduction as per Eqn.2.37.
In summary, Eqn.2.33 and 2.37 are conclusive as a proof of Prop.3 in the context of option PV and CVA, respectively.
2.2.4 Model Settings
Before proceeding with implementation details, we now introduce generic model settings for the underlying diffusion process. We consider multi-variate local volatility and single-asset stochastic volatility, both of which are representative examples beyond 1D by extending the dimension from multi-asset and multi-factor perspectives, respectively.
Multi-variate Local Volatility
for $i=1\ldots d$, where $q_{i}$ and $\sigma_{i}(S_{i}(t),t)$ are the constant continuous dividend yield and local volatility of the $i$-th asset’s motion, respectively; and $\{B_{i}(t)|i=1\ldots d\}$ is a correlated $d$-dimensional Brownian motion, with $\operatorname{\mathbb{E}}[dB_{i}(t)dB_{j}(t)]=\rho_{ij}dt:=[\delta_{ij}+\rho(1-\delta_{ij})]dt$ using Kronecker delta notation $\delta_{ij}$ and $\rho$ denoting single correlation between assets. Since we only consider continuous, rather than discrete, dividends, we assume $\vec{S}(0)=(1,\ldots,1)^{T}$ for simplicity without loss of generality. In the limit of $\sigma_{i}(S_{i}(t),t)$ being constants, this setting degenerates to the celebrated Black-Scholes model.
Stochastic Volatility: Heston Model
For stochastic volatility setting, let’s consider below stochastic differential equation under Heston model
with $\operatorname{\mathbb{E}}[dB^{S}(t)dB^{\nu}(t)]=\rho^{S,\nu}dt$, where $q$ is the constant continuous dividend yield, $\theta$ is the long variance, $\kappa$ is the mean reverting rate, $\xi$ is the volatility of volatility and $\rho^{S,\nu}$ is the correlation between spot and variance. For the same reason as previous subsection, we assume $S(0)=1$ for simplicity without loss of generality. And the initial variance is given by $\nu(0)$. There is a well known result on the expectation value of $\nu(t)$ over time: $\operatorname{\mathbb{E}}[\nu(t)]=\nu(0)e^{-\kappa t}+\theta(1-e^{-\kappa t})$.

### v8

@@ 2.3.3 Error Analysis
$e^{PV,P}_{k}\leq|\operatorname{\mathbb{E}}(B^{P}_{k})|+\mathcal{Y}^{P}_{k}e^{PV,P}_{k+1},$ (2.49)
where the identities of $B^{P}_{k}$ from Eqn.2.45 and $\mathcal{Y}^{P}_{k}$ from Eqn.2.46 have been applied, respectively; and $\leq$’ stems from a different sign of the covariance term compared to $\operatorname{\mathbb{E}}(V_{k}^{P}-V_{k})$ in Eqn.2.48, as analysed previously.
For $|\operatorname{\mathbb{E}}(B^{P}_{k})|^{2}$, we have,
$|\operatorname{\mathbb{E}}(B^{P}_{k})|^{2}$ $=$ $\operatorname{\mathbb{E}}((B^{P}_{k})^{2})-\operatorname{\mathrm{Var}}(B^{P}_{k})$
$=$ $\operatorname{\mathbb{E}}(|Z_{k}-F_{k}|^{2}|\Theta(F_{k}-Z_{k})\Theta(Z_{k}-P_{k}[V^{P}_{k+1}])-\Theta(P_{k}[V^{P}_{k+1}]-Z_{k})\Theta(Z_{k}-F_{k})|^{2})-\operatorname{\mathrm{Var}}(B^{P}_{k})$
$\leq$ $\operatorname{\mathbb{E}}(|Z_{k}-F_{k}|^{2}\Theta(|P_{k}[V^{P}_{k+1}]-F_{k}|-|F_{k}-Z_{k}|))-\operatorname{\mathrm{Var}}(B^{P}_{k})$
$\leq$ $\operatorname{\mathbb{E}}(|P_{k}[V^{P}_{k+1}]-F_{k}|^{2})\operatorname{\mathbb{E}}(\Theta(|P_{k}[V^{P}_{k+1}]-F_{k}|-|F_{k}-Z_{k}|))-\operatorname{\mathrm{Var}}(B^{P}_{k})$
$=$ $\operatorname{\mathbb{E}}(|P_{k}[V^{P}_{k+1}]-F_{k}|^{2})\mathcal{X}^{P}_{k}-\operatorname{\mathrm{Var}}(B^{P}_{k}),$
where we have made use of $\mathcal{X}^{P}_{k}$ (which in practice is small as $P_{k}[V^{P}_{k+1}]$ is close to $F_{k}$) from Eqn.2.45 in the penultimate line.
@@
where in the last line, the last term is exactly the local regression error with upper bound $\Xi_{k}^{P}[V_{k+1}]$ due to the projection on $P_{k}[V_{k+1}]$, which doesn’t consider accumulated error yet (as opposed to $P_{k}[V^{P}_{k+1}]$). To get rid of the two offsetting variance terms within $|\operatorname{\mathbb{E}}(B^{P}_{k})|^{2}$, we need to make use of Assumption 1, i.e., $\operatorname{\mathrm{Var}}(B^{P}_{k})\geq\operatorname{\mathrm{Var}}(V^{P}_{k+1}-V_{k+1})\cdot\mathcal{X}^{P}_{k}$. Then we arrive at
$\begin{split}|\operatorname{\mathbb{E}}(B^{P}_{k})|^{2}&\leq\mathcal{X}^{P}_{k}\cdot((e^{PV,P}_{k+1})^{2}+\Xi_{k}^{P}[V_{k+1}]+\operatorname{\mathrm{Var}}(V^{P}_{k+1}-V_{k+1}))-\operatorname{\mathrm{Var}}(B^{P}_{k})\\ &\leq\mathcal{X}^{P}_{k}\cdot((e^{PV,P}_{k+1})^{2}+\Xi_{k}^{P}[V_{k+1}]).\end{split}$ (2.50)
Note that as $V^{P}_{k+1}$ approaches the exact solution, both sides of the inequality in Assumption 1 tend to zero, it conjectures that the right-hand side, with $V^{P}_{k}-V_{k}\to 0$ and $\mathcal{X}^{P}_{k}\to 0$, converges faster than the left-hand side with finite $Z_{k}-F_{k}$. In practical terms, this assumption can be relaxed as
$\mathcal{X}^{P}_{k}\cdot(e^{PV,P}_{k+1})^{2}+\operatorname{\mathrm{Var}}(B^{P}_{k})\gg\mathcal{X}^{P}_{k}\cdot\operatorname{\mathrm{Var}}(V^{P}_{k+1}-V_{k+1}),$
in such a way that $\mathcal{X}^{P}_{k}\cdot(e^{PV,P}_{k+1})^{2}$ outweighs the difference between the two variance terms, even if that difference is positive.
Next, $\Xi_{k}^{P}[V_{k+1}]$ can be bounded by $\Xi_{Max}^{P}$ (for both LSM and FD-LSM, making use of Eqn.2.44 from Lemma 3) as
$\begin{split}\Xi_{k}^{P^{\text{LSM}}}[V_{k+1}]&\leq\Xi_{n^{*}}^{P^{\text{LSM}}}[V_{n^{*}+1}]\\ &=\Xi_{Max}^{P^{\text{LSM}}},\\ \Xi_{k}^{P^{\text{FD-LSM}}}[V_{k+1}]&=\frac{R-1}{R}[1-\tilde{\rho}_{k}^{2}(V_{k+1},\hat{V}^{1D}_{k+1})]\cdot\Xi_{k}^{P^{\text{LSM}}}[V_{k+1}]\\ &\leq\frac{R-1}{R}[1-\tilde{\rho}_{m^{*}}^{2}(V_{{m^{*}}+1},\hat{V}^{1D}_{{m^{*}}+1})]\cdot\Xi_{k}^{P^{\text{LSM}}}[V_{k+1}]\\ &\leq\frac{R-1}{R}[1-\tilde{\rho}_{m^{*}}^{2}(V_{{m^{*}}+1},\hat{V}^{1D}_{{m^{*}}+1})]\cdot\Xi_{n^{*}}^{P^{\text{LSM}}}[V_{n^{*}+1}]\\ &=\Xi_{Max}^{P^{\text{FD-LSM}}},\end{split}$ (2.51)
where $n^{*}=\arg_{n\in[1,M]}\max\Xi_{n}^{P^{\text{LSM}}}[V_{n+1}]$.
@@
Combining above with Eqn.2.49 and 2.50, the dynamical programming solution on the upper bound of $e^{PV,P}_{k}$, denoted by $\epsilon^{PV,P}_{k}$ for a given projection $P$, is
$\epsilon^{PV,P}_{k}=\sqrt{\mathcal{X}^{P}_{k}\cdot((\epsilon^{PV,P}_{k+1})^{2}+\Xi_{Max}^{P})}+\mathcal{Y}^{P}_{k}\epsilon^{PV,P}_{k+1},$ (2.52)
@@
with initial condition $\epsilon^{PV,P}_{M}=0$. It defines the relation between accumulated errors and local regression errors on option prices under least-squares Monte Carlo at each step.
The problem is now reduced to analyzing the properties of an iterative backward sequence defined by $x_{k}=\sqrt{a_{k}(x_{k+1}^{2}+b)}+c_{k}x_{k+1}$, where $a_{k}=\mathcal{X}^{P}_{k},b=\Xi_{Max}^{P},c_{k}=\mathcal{Y}^{P}_{k}$, with the initial condition $x_{M}=0$.
For the general multi-period setting under finite $M$, following Eqn.4.7 as a proof by induction in Appendix A.6.1, we have
$\epsilon^{PV,P^{\text{FD-LSM}}}_{k}\sim O(\sqrt{\Xi_{Max}^{P^{\text{FD-LSM}}}})\sim O(\sqrt{1-\tilde{\rho}_{m^{*}}(V_{m^{*}+1},\hat{V}^{1D}_{m^{*}+1})})$ (2.53)
as $\tilde{\rho}_{m^{*}}(V_{m^{*}+1},\hat{V}^{1D}_{m^{*}+1})\to 1$, whereas $\epsilon^{PV,P^{\text{LSM}}}_{k}\sim O(\sqrt{\Xi_{Max}^{P^{\text{LSM}}}})$ remains finite, for any $k\in[1,M]$.
Next, we consider the asymptotic behavior of the error in the limit of $M\to\infty$ and $k\to 1$. As we can see, if Assumption 2 also holds, the fixed point condition of Eqn.4.10 in Appendix A.6.2 is satisfied with $a_{0}=\mathcal{X}^{P}_{0}$ and $c_{0}=\mathcal{Y}^{P}_{0}$. Consequently, using Eqn.4.9 and substituting $b=\Xi_{Max}^{P}$, one can see that the error saturates at
$\epsilon^{PV,P}_{0}:=\lim_{\begin{subarray}{c}k\to 1\\ M\to\infty\end{subarray}}\epsilon^{PV,P}_{k}=\sqrt{\frac{\Xi_{Max}^{P}}{\frac{(1-\mathcal{Y}^{P}_{0})^{2}}{\mathcal{X}^{P}_{0}}-1}},$ (2.54)
@@
as going backward. As usual, we have $\epsilon^{PV,P^{\text{FD-LSM}}}_{0}\sim O(\sqrt{1-\tilde{\rho}_{m^{*}}(V_{m^{*}+1},\hat{V}^{1D}_{m^{*}+1})})$ in the limit of $M\to\infty$.
For a more tractable analysis of Eqn.2.52, let’s assume $\mathcal{X}^{P}_{k}=\mathcal{X}$ and $\mathcal{Y}^{P}_{k}=\mathcal{Y}$ for all $k$ in both LSM and FD-LSM.
Using $\tilde{\rho}_{m^{*}}(V_{m^{*}+1},\hat{V}^{1D}_{m^{*}+1})=90\%$ and $R=4$, $\Xi_{Max}^{P^{\text{LSM}}}=5\%^{2}$ and thus
$\Xi_{Max}^{P^{\text{FD-LSM}}}=\left(5\%\times\sqrt{\frac{4-1}{4}(1-90\%^{2})}\right)^{2}=1.89\%^{2}$ as a numerical example along with $\mathcal{X}=0.01$ and $\mathcal{Y}=0.8$, we have the plots for $\epsilon^{PV,P}_{k}$ as depicted in Fig.2.1. The asymptotic limits obtained through backward iteration are consistent with the theoretical predictions of Eqn.2.54 2.54: specifically, $\epsilon^{PV,P^{\text{LSM}}}_{0}=2.89\%$ for LSM and $\epsilon^{PV,P^{\text{FD-LSM}}}_{0}=1.09\%$ for FD-LSM.
Figure 2.1: The upper bound of accumulated option price error for LSM and FD-LSM at each step going backward, using the parameters mentioned in above. The $x$-axis is $M-k$ where M=60.
@@
Our quantitative result in Eqn.2.54 is consistent with previous qualitative study on the same matter by Stentoft, (2014): due to the fact that the projection approximation in LSM is only utilized to estimate stopping time rather than value function, the magnitude of accumulated error is relatively small. In fact, as stated by Tsitsiklis and Van Roy, (2001) in their Theorem 1, the accumulated error (upper bound) on option value scales as $\sqrt{M}$ when $k\to 1$ by applying the projection approximation to the value function itself. Now with a customized ansatz strongly correlated to the original problem, we show that theoretically in Eqn.2.53, FD-LSM could achieve even much lower level of accumulated errors.
Error analysis in CVA: Multi-bucket setting
@@
Now we turn to the error analysis of the CVA problem in a similar manner. As one could expect, the derivation of the proof for CVA problem in Eqn.2.15 is much less involved, due to the absence of a dynamic programming control coupled with least-squares projection, i.e., the local regression error due to $\tilde{F}_{k}\to\tilde{F}^{P}_{k}$ at each time step is decoupled and independent from other steps without being accumulated.
@@
by differentiating between $\Phi^{CVA}_{k}$ in Eqn.2.16 and its proxy defined as
@@
where $h_{cpty}(\tilde{F}^{P}_{k},t_{k})$ is approximated due to its dependence on $\tilde{F}^{P}_{k}$ subjected to the least-squares approximation.
Given the explicit functional dependence on $\tilde{F}_{i}$ ($i\in[1,k]$) in Eqn.2.56, the upper bound of $e^{CVA,P}_{k}$, denoted by $\epsilon^{CVA,P}_{k}$, can be derived as (defining $\Xi^{P}_{i}[\tilde{V}]:=\operatorname{\mathbb{E}}[(\tilde{F}^{P}_{i}-\tilde{F}_{i})^{2}]$)
$\begin{split}e^{CVA,P}_{k}&\leq\operatorname{\mathbb{E}}(|\Phi^{CVA}_{k}-\Phi^{CVA,P}_{k}|)\\ &\approx\operatorname{\mathbb{E}}[|\sum^{k}_{i=1}\frac{\partial\Phi^{CVA}_{k}}{\partial\tilde{F}_{i}}\cdot(\tilde{F}^{P}_{i}-\tilde{F}_{i})|]\\ &\leq\sum^{k}_{i=1}\operatorname{\mathbb{E}}[|\frac{\partial\Phi^{CVA}_{k}}{\partial\tilde{F}_{i}}\cdot(\tilde{F}^{P}_{i}-\tilde{F}_{i})|]\\ &\leq\sum^{k}_{i=1}\sqrt{\operatorname{\mathbb{E}}[\left(\frac{\partial\Phi^{CVA}_{k}}{\partial\tilde{F}_{i}}\right)^{2}]}\cdot\sqrt{\operatorname{\mathbb{E}}[(\tilde{F}^{P}_{i}-\tilde{F}_{i})^{2}]},\\ \epsilon^{CVA,P}_{k}&:=\sum^{k}_{i=1}\sqrt{\operatorname{\mathbb{E}}[\left(\frac{\partial\Phi^{CVA}_{k}}{\partial\tilde{F}_{i}}\right)^{2}]}\cdot\sqrt{\Xi^{P}_{i}[\tilde{V}]}\end{split}$ (2.57)
@@
To proceed, let’s denote $\tilde{m}^{*}=\arg_{m\in[1,M]}\min\tilde{\rho}_{m}(\tilde{V},\hat{\tilde{V}}^{1D})$, where $\tilde{\rho}_{k}(\tilde{V},\hat{\tilde{V}}^{1D})$ is the correlation between European options $\tilde{V}$ and $\hat{\tilde{V}}^{1D}$, a direct analogue of $\tilde{\rho}_{k}(V_{k+1},\hat{V}_{k+1}^{1D})$ in the previous problem. After some simple algebra and making use of Eqn.2.44, we have
$\begin{split}\epsilon^{CVA,P^{\text{FD-LSM}}}_{k}&=\sum^{k}_{i=1}\sqrt{\operatorname{\mathbb{E}}[\left(\frac{\partial\Phi^{CVA}_{k}}{\partial\tilde{F}_{i}}\right)^{2}]}\cdot\sqrt{\frac{R-1}{R}[1-\tilde{\rho}^{2}_{i}(\tilde{V},\hat{\tilde{V}}^{1D})]}\cdot\sqrt{\Xi^{P^{\text{LSM}}}_{i}[\tilde{V}]}\\ &\leq\sum^{k}_{i=1}\sqrt{\operatorname{\mathbb{E}}[\left(\frac{\partial\Phi^{CVA}_{k}}{\partial\tilde{F}_{i}}\right)^{2}]}\cdot\sqrt{\frac{R-1}{R}[1-\tilde{\rho}^{2}_{\tilde{m}^{*}}(\tilde{V},\hat{\tilde{V}}^{1D})]}\cdot\sqrt{\Xi^{P^{\text{LSM}}}_{i}[\tilde{V}]}\\ &=\sqrt{\frac{R-1}{R}[1-\tilde{\rho}^{2}_{\tilde{m}^{*}}(\tilde{V},\hat{\tilde{V}}^{1D})]}\epsilon^{CVA,P^{\text{LSM}}}_{k}.\end{split}$ (2.58)
It is worth to note that the CVA error under LSM could be difficult to control, since value function is approximated directly; in the new method, by virtual of the 1D ansatz as a control variate, it provides an invaluable error reduction $\epsilon^{CVA,P^{\text{FD-LSM}}}_{k}\to 0$ since $\tilde{\rho}_{\tilde{m}^{*}}(\tilde{V},\hat{\tilde{V}}^{1D})\to 1$, based on above Eqn.2.58.
In summary, Eqn.2.53 and 2.58 are conclusive as a proof of Prop.2 in the context of option PV and CVA, respectively.